# Low Resource Video Super-resolution using Memory and Residual Deformable Convolutions

### v1

@@ Abstract
Transformer-based video super-resolution (VSR) models have set new benchmarks in recent years, but their substantial computational demands make most of them unsuitable for deployment on resource-constrained devices. Achieving a balance between model complexity and output quality remains a formidable challenge in VSR. Although lightweight models have been introduced to address this issue, they often struggle to deliver state-of-the-art performance. We propose a novel lightweight, parameter-efficient deep residual deformable convolution network for VSR. Unlike prior methods, our model enhances feature utilization through residual connections and employs deformable convolution for precise frame alignment, addressing motion dynamics effectively. Furthermore, we introduce a single memory tensor to capture information accrued from the past frames and improve motion estimation across frames. This design enables an efficient balance between computational cost and reconstruction quality. With just 2.3 million parameters, our model achieves state-of-the-art SSIM of 0.9175 on the REDS4 dataset, surpassing existing lightweight and many heavy models in both accuracy and resource efficiency. Architectural insights from our model pave the way for real-time VSR on streaming data.
@@ 1 Introduction
Video Super-Resolution (VSR) is a fundamental task in computer vision that aims to reconstruct high-resolution videos from their low-resolution counterparts. This task is crucial in a wide range of applications, including video streaming, gaming, mobile content delivery, and surveillance. By leveraging both spatial and temporal information, VSR enhances the visual quality of videos while maintaining temporal consistency. However, achieving high-quality super-resolution under computational constraints remains a persistent challenge, particularly in real-world scenarios where resource efficiency is paramount.
@@
In recent years, deep learning-based approaches have shown remarkable progress in addressing the challenges of VSR. Convolutional architectures, such as EDVR and BasicVSR++, have achieved notable success by employing deformable convolutions to handle complex motion and residual networks to recover fine details. Transformer-based models, including VRT and RVRT, have further pushed the boundaries of VSR by capturing global temporal dependencies and improving long-term consistency. However, their substantial parameter count and computational requirements limit their scalability for lightweight applications.
@@
Recent efforts to design lightweight VSR models have explored motion-aware architectures and methods tailored for compressed video super-resolution. These approaches aim to balance performance and efficiency but often struggle to maintain temporal coherence or handle diverse real-world degradations. Furthermore, most existing methods focus solely on spatial and temporal domains, neglecting the potential benefits of incorporating frequency-domain representations for multi-scale feature extraction. This gap highlights the need for an innovative design that integrates these domains while maintaining computational efficiency.
Figure 1: Performance comparison of various Video Super-Resolution (VSR) models in terms of SSIM on the REDS4 dataset versus model parameters. The size of each bubble represents the computational complexity (GFLOPs) of the corresponding model. Our proposed method, RCDM, achieves a competitive SSIM score with significantly lower computational cost compared to transformer-based methods such as VRT, PSRT, and RVRT.
@@
In this work, we propose a novel lightweight VSR framework that combines convolutional efficiency with advanced spatiotemporal modeling to achieve state-of-the-art results under resource constraints. Our architecture builds upon the strengths of convolutional networks, which excel at localized feature extraction, and enhances them with components tailored for VSR-specific challenges. By integrating deformable convolutions, we ensure effective motion alignment across frames, which is crucial for maintaining temporal consistency. To improve spatial and structural representation, we introduce a dual-pipeline design that incorporates 2D Discrete Wavelet Transform (DWT) for frequency-domain feature extraction. This approach not only captures fine details but also reduces computational overhead by focusing on multi-scale representations. Furthermore, we leverage memory tensors to enhance temporal coherence, enabling our model to handle dynamic and complex video scenarios more effectively. Additionally, ConvNext blocks are employed to ensure efficient spatial feature extraction while maintaining a low parameter count and GFLOPs.
@@ 2 Related Work
Video Super-Resolution (VSR) Overview: Video super-resolution has been extensively studied, with recent advancements primarily focusing on leveraging deep learning models to enhance resolution by utilizing spatial and temporal dependencies across consecutive frames. While early methods relied on explicit motion compensation, modern approaches employ deeper architectures and attention mechanisms to model these dependencies implicitly or explicitly.
Transformer-Based VSR Models: Transformer-based architectures, such as VRT [3] and RVRT [4], have achieved impressive performance in VSR tasks by utilizing global attention mechanisms to model long-term temporal dependencies. However, these methods often come with substantial computational overhead, making them less suitable for real-time or resource-constrained scenarios. While transformers are highly effective in global context modeling, their high parameter count and memory requirements hinder practical deployment, particularly in applications requiring low latency [5].
Lightweight VSR Models: With the increasing demand for computational efficiency, lightweight VSR models have been explored extensively. ELNVSR [6] proposed a bidirectional alignment module and multi-scale pyramid block, effectively capturing spatial-temporal correlations while reducing complexity. Similarly, IRN [7] utilized ConvNeXt blocks and residual structures to balance performance and computational cost. Although these methods reduce model size and computation, they often underperform in frame alignment and temporal consistency, particularly in dynamic scenes.
Figure 2 illustrates the network pipeline for our family of architectures.
Deformable Convolutions in VSR: Deformable convolutional layers, as introduced in models like EDVR [2] and BasicVSR++ [1], have become a cornerstone of VSR for handling motion misalignment. These layers dynamically predict spatial offsets, enabling robust motion alignment across frames, even in challenging scenarios with large displacements. Despite their effectiveness, the integration of deformable convolutions increases model complexity, necessitating efficient implementations for lightweight applications [6].
Memory-Augmented Architectures: Memory mechanisms have been proposed to enhance temporal coherence in VSR models by retaining information from previous frames. RealBasicVSR [8] introduced a dynamic refinement module that iteratively reduces artifacts during propagation. By leveraging memory for long-term dependency modeling, these architectures achieve smoother transitions between frames. However, their reliance on large memory buffers can pose challenges for resource-limited settings.
@@
Efficient Upsampling and Wavelet-Based Methods: Efficient upsampling techniques have also been explored to reduce computational overhead. Dynamic Context-Guided Upsampling (DCGU) [9] demonstrated effective texture enhancement and scale adaptation by leveraging non-local sampling and filter weight modulation. Additionally, wavelet-based methods have gained attention for their ability to capture multi-scale frequency information. These approaches provide an efficient mechanism for structural preservation and fine detail reconstruction, making them well-suited for lightweight VSR architectures.
Real-World VSR Challenges and Solutions: Models designed for real-world VSR, such as RealBasicVSR [8], emphasize balancing detail synthesis and artifact suppression. These methods address the challenges of unknown degradations by integrating pre-cleaning stages and dynamic refinement schemes. Real-world datasets like RealVSR [10] and VideoLQ [8] have further facilitated benchmarking and model evaluation, providing a foundation for developing robust architectures.
Our Contributions: Building upon the advancements in deformable convolutions, memory mechanisms, and lightweight architectures, we propose a novel convolution-based VSR model tailored for resource-constrained settings. Unlike transformer-based methods, our approach prioritizes computational efficiency through the integration of residual networks, deformable convolutions, wavelet transforms, and memory tensors. By leveraging these components, our model achieves a balance between performance and complexity, addressing key challenges in frame alignment, temporal consistency, and structural preservation. Comprehensive comparisons with state-of-the-art methods validate the effectiveness of our approach in delivering high-quality video reconstruction with minimal computational requirements.
@@ 3.1 Problem Formulation
Given a low-resolution (LR) video sequence $\{I_{t-N},\dots,I_{t},\dots,I_{t+N}\}$, where $I_{t}$ represents the $t$-th frame and $N$ defines the temporal radius, our goal is to reconstruct a high-resolution (HR) frame $H_{t}$ corresponding to $I_{t}$. The VSR process can be formulated as a mapping function:
$H_{t}=f(I_{t-N},\dots,I_{t},\dots,I_{t+N};\theta),$ (1)
where $f(\cdot)$ is a deep neural network parameterized by $\theta$, responsible for restoring missing high-frequency details while preserving temporal consistency.
Mathematical Interpretation of VSR
VSR involves restoring high-resolution details from a set of degraded low-resolution frames. The degradation process can be modeled as:
@@
where $B(\cdot)$ represents blur, $D(\cdot)$ denotes downsampling, and $\eta_{t}$ models additive noise. The primary goal of our network is to approximate the inverse mapping:
$H_{t}\approx f_{\theta}(I_{t-N},...,I_{t},...,I_{t+N}),$ (3)
where the function $f_{\theta}$ learns to recover fine details lost in $B(\cdot)$ and $D(\cdot)$ while maintaining motion alignment across frames.
3.2 Replacing Transformers with ConvNeXt Blocks
Transformers have shown remarkable performance in video super-resolution tasks due to their ability to model long-range dependencies. However, their quadratic complexity in self-attention makes them inefficient for real-time applications. Instead of adopting transformer-based architectures like VRT [3] and RVRT [4], we employ ConvNeXt [11], a modernized convolutional architecture that integrates design principles from transformers while retaining the computational efficiency of ConvNets.
ConvNeXt layers are defined as:
$y=W_{2}\cdot\text{GELU}(W_{1}\cdot\text{DWConv}(x))+x,$ (4)
where $x$ represents the input feature map, $\text{DWConv}(\cdot)$ is a depthwise convolution, and $W_{1},W_{2}$ are learnable weight matrices. The use of GELU non-linearity and residual connections improves expressivity while maintaining efficiency.
3.3 Network Architecture
Our proposed architecture follows a convolution-based approach that integrates deformable convolutions, residual connections, wavelet transforms, and memory propagation mechanisms to enhance video resolution efficiently. Figure 2 illustrates the network pipeline.
Figure 2: Proposed low-resource VSR model. The network extracts multi-frame features using deformable convolutions, wavelet transforms, and memory tensors, ensuring both spatial and temporal consistency.
3.4 Deformable Motion Compensation
Traditional VSR models rely on optical flow for frame alignment, which introduces artifacts when dealing with complex motion. Instead, we employ Deformable Conv3D, which dynamically adapts the receptive field to handle motion across frames:
$F_{t}=\sum_{i=1}^{K}w_{i}\cdot I_{t+\Delta t_{i}},$ (5)
where $w_{i}$ represents the learned spatial offsets and $\Delta t_{i}$ indicates the adaptive displacement across adjacent frames. This formulation ensures that motion information is effectively captured without explicit optical flow estimation.
@@ 3.5 Wavelet-Based Multi-Scale Feature Extraction
To enhance texture details and reduce redundancy, we integrate 2D Discrete Wavelet Transform (DWT) into our architecture. The DWT decomposition of the central frame is given by:
$I_{t}^{LL},I_{t}^{LH},I_{t}^{HL},I_{t}^{HH}=\text{DWT}(I_{t}),$ (6)
where $LL,LH,HL,HH$ represent the approximation, horizontal, vertical, and diagonal detail coefficients, respectively. These coefficients are processed separately and later fused to reconstruct the HR frame, providing improved structural integrity.
3.6 Memory Tensor for Temporal Propagation
To improve temporal consistency, we introduce a memory tensor module that stores historical feature representations. Instead of relying solely on recurrent structures, we use a residual memory update mechanism:
$F_{t}=\alpha\cdot M_{t-1}+\cdot F_{t},$ (7)
where $\alpha$ is a learnable weighting parameter that determines the contribution of past information. This memory mechanism ensures long-term temporal coherence, preventing flickering artifacts in reconstructed videos. Here $\ F_{t}$ represents the current super resolved frame while $\ M_{t-1}$ represents the memory tensor generated from the previous frame.
3.7 Upsampling and Reconstruction
After extracting and fusing spatial-temporal features, the final HR frame $H_{t}$ is reconstructed using an efficient upsampling module. We utilize a pixel-shuffle layer to upscale the feature maps:
$H_{t}=\sigma(W\ast P(U_{t})),$ (8)
where $P(U_{t})$ represents the pixel-shuffle operation, $W$ is the learnable weight matrix, and $\sigma(\cdot)$ is the Sigmoid activation function ensuring pixel intensity normalization.
3.8 Performance and Computational Trade-offs
Replacing transformers with ConvNeXt blocks significantly reduces computational overhead while maintaining competitive performance. Table 1 compares different model variants.

### v2

# Low-Resource Video Super-Resolution using Memory, Wavelets, and Deformable Convolutions

@@ Abstract
The tradeoff between reconstruction quality and compute required for video super-resolution (VSR) remains a formidable challenge in its adoption for deployment on resource-constrained edge devices. While transformer-based VSR models have set new benchmarks for reconstruction quality in recent years, these require substantial computational resources. On the other hand, lightweight models that have been introduced even recently struggle to deliver state-of-the-art reconstruction. We propose a novel lightweight and parameter-efficient neural architecture for VSR that achieves state-of-the-art reconstruction accuracy with just 2.3 million parameters. Our model enhances information utilization based on several architectural attributes. Firstly, it uses 2D wavelet decompositions strategically interlayered with learnable convolutional layers to utilize the inductive prior of spatial sparsity of edges in visual data. Secondly, it uses a single memory tensor to capture inter-frame temporal information while avoiding the computational cost of previous memory-based schemes. Thirdly, it uses residual deformable convolutions for implicit inter-frame object alignment that improve upon deformable convolutions by enhancing spatial information in inter-frame feature differences. Architectural insights from our model can pave the way for real-time VSR on the edge, such as display devices for streaming data.
@@ 1 Introduction
Video Super-Resolution (VSR) aims to reconstruct high-resolution videos from their low-resolution counterparts. This task is crucial in a wide range of applications, including video streaming, gaming, mobile content delivery, and surveillance. By leveraging both spatial and temporal information, VSR enhances the visual quality of videos while maintaining temporal consistency. However, achieving high-quality super-resolution under computational constraints remains a persistent challenge, particularly in scenarios where resource efficiency is paramount.
Figure 1: Comparison of video super-resolution (VSR) models: Proposed models (blue filled circles) reconstruct frames with competitive quality (SSIM on REDS4) when compared to previous models (grey hollow circles) with fewer trainable parameters and inference floating point operations per frames (GFLOPs in parenthesis for model labels and also coded as bubble size).
@@
In recent years, deep learning-based approaches have shown remarkable progress in addressing the challenges of VSR. Convolutional architectures, such as video restoration framework with enhanced deformable networks (EDVR) [25] and BasicVSR++ [2], have achieved notable success by employing deformable convolutions and residual networks to utilize motion information and recover fine details, respectively. Transformer-based models, including video restoration transformer (VRT) and recurrent video restoration transformer (RVRT), have further pushed the boundaries of VSR by capturing global temporal dependencies and improving long-term consistency[13][14]. However, their substantial parameter count and computational requirements limit their utility for deployment in setups with strict power, compute, and cost constraints.
@@
VSR models have explored motion-aware architectures such as VESPCN[1] and EDVR[25], as well as methods tailored for compressed video super-resolution[8]. These approaches aim to balance performance and efficiency, yet they often struggle with temporal coherence, handling large motion, and adapting to diverse real-world degradations—particularly in highly compressed or low-quality videos with unknown distortions.Furthermore, most existing methods focus solely on spatial and temporal domains, neglecting the potential benefits of incorporating frequency-domain representations for multi-scale feature extraction. This gap highlights the need for an innovative design that integrates these domains while maintaining computational efficiency.
@@
In this work, we propose a novel lightweight VSR framework that combines convolutional efficiency with advanced spatiotemporal modeling to achieve state-of-the-art results with limited resources. Our architecture builds upon the strengths of convolutional networks, which excel at localized feature extraction, and enhances them with components tailored for challenges specific to VSR. To improve spatial and structural representation, we introduce a dual-pipeline design that incorporates 2D discrete wavelet transform (DWT) for multi-resolution and spatially sparse edge feature extraction and 3D convolutional residual blocks for spatio-temporal feature extraction. Additionally, we integrate deformable convolutions to ensure effective motion alignment across frames, which is crucial for maintaining temporal consistency. Furthermore, we leverage a memory tensor to enhance utilization and reconstruction of temporal coherence. Additionally, ConvNeXt blocks are employed to ensure efficient spatial feature extraction while maintaining a low parameter count and GFLOPs.
@@
Following are our key contributions:
  • State-of-the-art reconstruction quality on REDS4[21]: Our method achieves an SSIM of 0.9175, surpassing much heavier models, such as BasicVSR++[2] (0.9069) and RVRT (0.9113)[14].
  • Low computational overhead compared to previous models: Transformer-based models such as VRT (30.7M parameters)[13] and RVRT (11.06M parameters)[14] exhibit high computational costs, while our model delivers comparable or better performance at significantly lower resource requirements. The floating point operations required at test time are also from seven to 35 times lower than comparable models (see Figure 1).
  • Memory tensor enhancements: Our introduction of memory tensors plays a key role in preserving temporal consistency and significantly improving SSIM scores across multiple datasets. Compared to previous memory-based approaches, such as those using convolutional LSTM, our approach is significantly more compute efficient. This approach enhances temporal stability without excessive model complexity.
  • Residual deformable convolutions: Our work builds on the strengths of deformable convolutions for motion alignment but incorporates residual deformable convolutions, which improve feature reuse and reduce the need for excessive parameterization.
  • Wavelet decomposition in 2D: We utilize the multi-resolution and energy compaction (sparsification) properties of 2D wavelets when applied to natural images to increase the spatial information utilization in our architecture.
  • Orchestration of multiple architectural features: While ideas such as 3D convolutions and residual blocks have been used in VSR before, our orchestration of multiple parallel branches that specialize in exploiting various aspects of spatio-temporal information available in videos is unique, which reflects in its results.
@@ 2 Related Work
Transformer-Based Video Super-Resolution (VSR) Models: Transformer-based architectures, such as Video Restoration Transformer (VRT) citeliang2022vrt and Recurrent Video Restoration Transformer (RVRT)[14], leverage global attention to model long-term temporal dependencies, achieving state-of-the-art performance in VSR. However, their high computational cost and memory footprint make them impractical for real-time applications or deployment on resource-constrained devices [16]. Progressive Sparse Representation Transformer (PSRT) [22] improves efficiency using a recurrent design but still suffers from high latency. While these transformer-based methods achieve strong reconstruction quality, their computational and memory requirements limit their utility for deployment in low-resource settings.
Lightweight VSR Models: Several recent works have focused on reducing model complexity while maintaining competitive performance. Efficient Lightweight Network for Video Super-Resolution (ELNVSR) [29] introduces a bidirectional alignment module and a multi-scale pyramid block to reduce redundancy and improve feature aggregation. Similarly, Information Refinement Network (IRN)[19] employs ConvNeXt blocks with residual structures, effectively balancing parameter efficiency and accuracy. However, these models still struggle with motion compensation and temporal coherence, particularly in challenging real-world scenarios.
A notable contribution in lightweight VSR is WaveMixSR[11], which employs wavelet transforms to extract multi-scale representations. By leveraging joint spatial and frequency-domain processing, WaveMixSR enhances texture restoration while reducing redundant spatial computations. However, despite its advantages, it does not integrate effective memory mechanisms for long-term temporal dependencies, leading to frame inconsistencies in videos.
Deformable Convolutions for Motion Alignment: Deformable convolutions have proven highly effective in handling motion misalignment in VSR. Enhanced Deformable Video Restoration (EDVR) [25] introduced deformable alignment layers that dynamically predict spatial offsets for better motion compensation. Basic Video Super-Resolution++ (BasicVSR++) [3] further refined this approach by incorporating bidirectional propagation with deformable convolutions, leading to improved performance in dynamic scenes. However, while these methods enhance motion alignment, they often introduce additional parameters, increasing computational complexity.
Memory-Augmented Architectures for Temporal Consistency: To address temporal inconsistencies in VSR, several models integrate memory-based architectures. RealBasicVSR [4] introduces a dynamic refinement module that progressively reduces artifacts by leveraging long-term dependencies. Similarly, RVRT [14] incorporates memory mechanisms to improve video consistency across frames. However, these methods still rely on high computational budgets, making them less viable for low-resource applications.
@@
Efficient Upsampling and Multi-Scale Feature Extraction: Efficient upsampling plays a crucial role in lightweight VSR models. Dynamic Context-Guided Upsampling (DCGU) [9] improves detail reconstruction by leveraging non-local sampling, enhancing fine-texture restoration with reduced computational cost. Meanwhile,wavelet-based methods have gained attention due to their ability to capture structural information across different frequency scales.
We integrate wavelet-based multi-scale feature extraction in our framework to improve detail preservation while reducing redundant computations. Unlike traditional spatial convolutions, wavelet transformations enable a more efficient and parameter-effective method of capturing high-frequency details.
Building upon these advancements, we propose a convolution-based VSR model that achieves state-of-the-art performance while remaining computationally efficient. Unlike transformer-heavy methods, our approach incorporates:
  • Residual Deformable Convolutions for precise frame alignment with minimal overhead.
  • Memory Tensors to enhance long-term temporal consistency without the need for excessive memory buffers.
  • Wavelet-Based Feature Extraction for multi-scale detail enhancement while reducing computational redundancy.
  • Parameter-Efficient Design, achieving superior performance on REDS4 and other benchmark datasets with only 2.3M parameters and significantly fewer FLOPs compared to transformer-based approaches.
Our model demonstrates that by strategically integrating memory mechanisms, residual networks, and deformable convolutions, it is possible to achieve high-quality VSR while maintaining low computational costs, making it suitable for real-world applications such as video streaming, surveillance, and mobile devices.
@@ 3 Proposed Method
We introduce Residual ConvNeXt Deformable Convolution with Memory (RCDM) – a VSR architcture that is designed for computational efficiency while maintaining state-of-the-art performance. Unlike transformer-based methods, RCDM employs deformable convolutions for motion compensation, wavelet-based feature extraction for texture enhancement, and memory tensors for long-term temporal consistency.
@@ 3.1 Problem formulation
Video super-resolution involves restoring high-resolution details from a set of degraded low-resolution frames. The degradation process can be modeled as:
@@
where $B(\cdot)$ represents blur, $D(\cdot)$ denotes downsampling, and $\eta_{t}$ models additive noise. The primary goal of our network is to approximate the inverse mapping.
Given a sequence of $12N+$ 1 LR frames $\{I_{t-N},...,I_{t},...,I_{t+N}\}$, the goal of RCDM is to reconstruct the corresponding central HR frame $H_{t}$. The LR frame is loaded to have the size of $(2N+1,b,c,h,w)$, which corresponds to frames, batch, channels, height, width while the generated HR has a size $(1,b,c,4h,4w)$ corresponding to the batch of central frames. Mathematically, this can be formulated as:
$H_{t},M_{t}=f_{\alpha}(I_{t-N},...,I_{t},...,I_{t+N},M_{t-1}),$ (2)
where $f_{\alpha}$ is the learned VSR model parameterized by $\alpha$, and $M_{t-1}$ is a memory tensor storing information from previous frames to ensure temporal consistency. The model outputs both the estimated HR frame $H_{t}$ and the memory tensor to be utilized in the next frame $M_{t}$.
3.2 Alignment Block
A key challenge in VSR is motion alignment across consecutive frames to account for movements of the camera or objects. Instead of relying on explicit optical flow estimation, which can introduce artifacts, we utilize Deformable Convolution for motion compensation to dynamically adapt the receptive field and align adjacent frames with the reference frame. A standard convolution samples feature values from a fixed grid, but a deformable convolution introduces learned offsets to adjust sampling locations dynamically[7]. Extending this to 3D allows the network to model both spatial and temporal motion, ensuring frame alignment without explicit motion estimation. The alignment block followed by feature extraction increases in channel depth while keeping the length and width as same as the LR frame (h,w)
3.3 Residual Spatio-Temporal Feature Extraction
Once aligned, the neighboring frame features are processed using a Residual 3D Convolution Block, which aggregates multi-frame information while preserving the structure of the reference frame. Residual connections extract spatio-temporal features, and ensures the preservation of base frame details.
@@ 3.4 Wavelet-Based Multi-Scale Feature Extraction
To enhance fine texture details, RCDM employs a 2D Discrete Wavelet Transform (DWT) to decompose the central frame into frequency sub-bands capturing horizontal, vertical, and diagonal details. These sub-bands are processed independently and later fused to reconstruct high-quality textures. Due to the DWT, the size of features changes from $(h,w,c)$ to $(h/2,w/2,4c)$ without any loss of information. To enable feature fusion by concatenation, an up-sampling block prior to fusion is processed.
3.5 Memory Mechanism for Temporal Consistency
To maintain long-term consistency across video frames, we introduce a memory tensor $M_{t}$, which accumulates information over time using a residual update mechanism:
$M_{t}=\beta M_{t-1}+H_{t}^{\text{feat}},$ (3)
where $\beta$ is a learnable weight controlling the influence of past frames, and $H_{t}^{\text{feat}}$ represents the refined features of the current frame. This allows the model to propagate previously learned details forward, preventing flickering artifacts in video sequences[5] [6].
3.6 Super-Resolution and Upsampling
The final stage refines the frame representation using ConvNeXt Blocks, which consist of depthwise convolutions and residual connections.[18] The ConvNeXt blocks enable feature refinement while keeping computational cost low.
Finally, upsampling is performed using a pixel shuffle layer[23], which efficiently rearranges feature maps to reconstruct the HR frame. In this block, HR upsampled versions are generated using pixel shuffle and ConvNeXT are combined for residual learning. This is then followed by learning from memory block to generate the expected final HR frame.
The complete RCDM pipeline is shown in Figure 2, which shows how it integrates:
  1. 1.
    Deformable Convolution for Motion Compensation: Aligns neighboring frames dynamically.
  2. 2.
    Residual Feature Fusion: Extracts spatio-temporal information while preserving details.
  3. 3.
    Wavelet-Based Processing: Enhances multi-scale textures.
  4. 4.
    Memory Tensor: Ensures temporal stability.
  5. 5.
    ConvNeXt Refinement and Upsampling: Produces high-quality HR frames.
At each timestep $t$, the model reconstructs $H_{t}$ while updating the memory for subsequent frames, ensuring smooth, artifact-free video super-resolution.
Figure 2: Overview of the proposed RCDM architecture. The model integrates deformable convolutions, wavelet decomposition, memory propagation, and ConvNeXt refinement for efficient video super-resolution.